# Progenitor diversity in the accreted stellar halos of Milky Way-like galaxies

### v1

@@ 2.3 Auriga
Auriga is a suite of N-body magnetohydrodynamical (MHD) zoom simulations of Milky Way-like galaxies from cosmological initial conditions (Marinacci & Pakmor, 2014; Grand et al., 2017, 2024). The simulations were carried out with the AREPO code (Springel, 2010). The original [^1] The recent Auriga data release described in Grand et al. (2024) includes 9 additional halos with lower virial masses, $10^{11}--10^{12}\,\mathrm{\mathrm{M_{\odot}}}$. These are not considered here or in earlier papers about stellar halos in Auriga. Auriga suite comprises 30 halos simulated at a baseline resolution (level 4, L4), labeled Au-1 to Au-30. Of these, 6 were re-simulated at even higher resolution (level 3, L3). We use only the L4 halos in this work. Our study is concerned with the bulk structure of the halos, which for the most part is dominated by the most massive accreted progenitors. The lower resolution of L4 is therefore not as significant as the benefit of a much larger sample.
@@
In contrast to Aquarius and BJ05, Auriga includes star formation and the dynamical evolution of gas and stars self-consistently. This is particularly important for modeling the growth of the central galaxy and its response to individual accretion events. The distribution of orbits for halo stars should therefore better reflect both the phase space distribution of stars in satellites before they are disrupted, and interactions between the central galaxy and the accreted satellites (including stronger dynamical friction, more rapid disruption and tidal shocking). In principle, a hydrodynamical treatment also allows for an ‘in situ halo’ formed by stars ‘kicked’ out of the disk during mergers or in gas clouds stripped from satellites, before they have time to mix into the gaseous halo or interstellar medium of the central galaxy (e.g. Cooper et al., 2015). However, as discussed by Cooper et al. (2017), it is not obvious that these effects will dominate the differences between particle tagging and hydrodynamical realizations of halos simulated from the same initial conditions: such differences are more likely to be dominated by model-to-model variations in the star formation histories of individual satellites predicted by different combinations of hydrodynamical and subgrid methods.
@@
The Auriga galaxies were chosen from a lower-resolution cosmological volume, based on a set of criteria to identify Milky Way analogues on the basis of their virial mass and isolation (see Grand et al., 2017, 2024). The resulting galaxies have radial scale lengths ranging from $2.16\,\mathrm{kpc}$ to $11.64\,\mathrm{kpc}$, most with prominent central bulge components. As described in Monachesi et al. (2019), their stellar halos, defined kinematically by stellar particles with circularity parameter $\epsilon<0.7$ and radius $>5$ kpc (to avoid the bulge region), have masses in the range $10^{9}$$10^{10}~{}\rm M_{\odot}$, with a significant contribution from stars formed in situ. The Auriga halos have a wide range of properties, consistent with the Aquarius halos in C10. Previous studies of Auriga have the above kinematic definition of the stellar halo, and other definitions; for consistency with the BJ05 and Aquarius models, we use a simple merger tree-based definition in this paper, described below.
@@ 2.4 Definition of the accreted stellar halo and its progenitors
Particles formed in long-lived satellites of the main halo may not be accreted into the stellar halo of the Milky Way analog until long after their parent halo has crossed the virial radius of the main branch (the infall time of the progenitor). A given progenitor ID therefore corresponds to a single infall time, but may be associated with a wide range of individual particle accretion times. Satellite halos may continue to form stars after they become subhalos; if such stars are unbound from their parent satellite at $z=0$, we label them as accreted, consistent with the treatment of stars formed in the same progenitor before it became a subhalo of the main branch.
@@
Note that the progenitor ID indicates the branch from which a star was unbound when it became part of the stellar halo of the Milky Way analogue: this is not necessarily the branch in which the star formed, because each subhalo joining the main branch has its own set of hierarchical progenitor branches. In the case of the most massive satellites, more than one of these ‘satellite of satellite’ branches may form stars. If a satellite-of-a-satellite is stripped from its host after infall, such that some of its stars are accreted directly into the stellar halo from the Milky Way analog, our algorithm will naturally assign a distinct progenitor ID to those stars. If those stars are unbound into their host before they join the main branch, they will instead be labelled with the progenitor ID of their host. Thus, stars formed in a single progenitor branch may have different progenitor IDs in some cases; this is consistent with our intended purpose here and is likely rare in practice [^3] Several technical subtleties are involved building merger trees by linking structures of halo-finding methods (like SUBFIND) across the snapshots of $N$-body simulations. For example, during complex merger events, the operational definition of the main branch as the most massive progenitor can be ambiguous, and links between ‘related’ branches or sections of branches can be broken (see Jiang et al., 2014, for extended discussion of the potential problems). The most significant impact of such uncertainties on our work is a small number of stellar particles that cannot be ascribed to any accretion event. We have not observed any cases in which these problematic particles contribute significant mass to the accreted stellar halo..
@@
In hydrodynamical simulations like Auriga, in situ stars can form directly on halo-like orbits or be scattered onto such orbits by gravitational interactions, after having formed in the central disk (e.g. Abadi et al., 2003; Zolotov et al., 2009; Cooper et al., 2015; Font et al., 2020). Alternative definitions of the stellar halo may count some fraction of these as “halo” stars in order to make a more realistic comparison with observations (for example, selecting stars on very eccentric orbits). We prefer an idealized separation of accreted and in situ stars, using knowledge of the merger history from the simulations, because our purpose is mainly to compare different models (Auriga may overpredict the mass of in situ stars on halo orbits, which complicates any comparison that mixes accreted and in situ halo stars Monachesi et al., 2019). In future work it would be useful to examine how in situ stars affect inferences about the composition of the accreted stellar halo.
@@ 2.5.1 Stellar halo masses
Figure 1: Virial masses and accreted stellar halo masses of galaxies in the three model sets we use in this work (symbols, as shown in the legend). The star symbol shows an approximation to currently favoured values for the Milky Way for reference ($M_{200}\sim 10^{12}\,\mathrm{M_{\odot}}$, $M_{\star}\sim 10^{9}\,\mathrm{M_{\odot}}$; e.g. Deason et al., 2019).
[…]
As we discuss below, despite this large system-to-system scatter, the diversity of individual stellar halos shows clear systematic differences between the three model sets. We note that Auriga provides the simulations that are closest to measurements of the virial mass and stellar halo mass of the Milky Way (e.g. Au09, Au10, Au17 and Au22).
@@ 2.5.2 Progenitor masses
Figure 2: The relation between stellar mass and dark matter halo mass (prior to accretion) for progenitor satellites in the 30 Auriga and 11 BJ05 models. Progenitors from all individual halos in each simulation set are combined in this plot. The top and right-hand panels show the normalized distributions of halo mass and stellar mass, respectively.
Fig. 2 shows the relationship between stellar mass and dark halo mass (approximately virial mass) for all progenitor satellites in Auriga (pink) and BJ05 (grey). The total virial mass here is defined prior to infall of the progenitor to the main halo, hence this plot shows the ‘underlying’ relation, without the effects of tidal stripping on surviving satellites. The figure provides some insight into the different star formation efficiencies for satellites predicted by the two models. They predict almost identical luminosity function shapes (righthand histogram), albeit with an excess of very low mass satellites in Auriga at masses comparable to the gas particle mass. Even well-resolved satellites in Auriga occupy somewhat less massive halos than satellites of comparable stellar mass in BJ05 (this can be seen in the shift of the halo mass distribution to lower halo mass for Auriga progenitors, as shown in the top histogram). The joint distributions show similar overall trends, but it is clear that Auriga progenitors have lower virial mass at fixed stellar mass (or equivalently, higher stellar mass at fixed virial mass). The difference is greatest (up to $\sim 1$ order of magnitude) for the most massive satellites.
The most significant consequence of these differences is that the most massive progenitors in Auriga can contribute relatively more mass to the accreted stellar halo. Together with other effects such as dynamical friction, this is likely to lead to the systematic differences between the diversity profiles of the two models, as we discuss further below.
@@ 2.5.3 Overall density profiles
Fig. 3 shows the overall distribution of accreted stars in the different models. The radial scales and outer power-law slopes are broadly consistent among the three models, but there are also clear systematic differences. The Auriga halos have the highest average density at all radii and are notably similar to one another in their outskirts, with $\lesssim 1~{}\mathrm{dex}$ variation in density between different galaxies at all radii $\gtrsim 10~{}\mathrm{kpc}$. The BJ05 halos have lower density than Auriga at radii $\lesssim 30~{}\mathrm{kpc}$ and even less galaxy-to-galaxy variation. The lower central densities of the BJ05 models (and perhaps also their lower average accreted stellar masses, at fixed virial mass) may be the result of the limitations of those models at higher redshift, as discussed above.
The Aquarius halos have lower density than most of the BJ05 and Auriga halos at radii $\gtrsim 10~{}\mathrm{kpc}$ and show greater variation from galaxy to galaxy, as well as a wider range of slopes and more pronounced inflections (‘breaks’). These breaks correspond to transitions between regions dominated by different combinations of progenitors, as described in C10. In some (not necessarily all) cases the breaks correspond to approximately to the apocentres of stars from individual dominant progenitors, and a pronounced break in the overall profile may indicate the mass of the accreted stellar halo overall is dominated by concentrated debris from a single progenitor (Deason et al., 2013). As we explore further below, the most significant difference between the Auriga and Aquarius stellar halos is that debris from individual massive satellites generally makes a dominant contribution to the overall density profile across a much wider range of radii in Auriga.
Figure 3: A comparison of the volume density profiles of the accreted stellar halos in each simulation. The 6 Aquarius halos are shown with individual colors. For simplicity, the 11 B & J halos and 30 Auriga halos are shown with gray and pink lines respectively.
@@ 3.1 Definition of diversity
Other measures of overall stellar halo diversity (i.e. for the whole accreted stellar halo) have been used in the literature. For example, as noted above, BJ05 reported the fraction of accreted stellar mass contributed by the 15 most massive progenitors. Deason et al. (2016) reported the mass-weighted average progenitor stellar mass, which (as they explain) is directly related to our $N_{\mathrm{sig}}$. Another example is that of Monachesi et al. (2019), who used a measure we call $N_{90}$, defined as the number of the progenitors required to account for 90% of the accreted stellar halo mass, when accumulated in descending order of progenitor mass and rounded up to the nearest integer (a similar measure is discussed in C10). For example, a halo in which the most massive progenitor accounts for $80\%$ of the accreted stellar halo mass, and the second most massive progenitor for a further $15\%$, would have $N_{90}=2$. By construction, $N_{\mathrm{90}}\geq N_{\mathrm{sig}}$. When a few massive progenitors account for most of the mass, but not more 90%, and the remaining mass up to 90% is dominated by a relatively large number of much less massive progenitors, then $N_{90}$ exceeds $N_{\mathrm{sig}}$ by a larger factor.
@@
Fig. 4 compares $N_{90}$ with $N_{\mathrm{sig}}$ for the halos in the three simulation sets. It is clear that the different models predict different distributions of progenitor mass for the accreted stellar halo. The BJ05 and Auriga models contain examples of systems in which most of the accreted mass is contributed by many progenitors, each only a small fraction of the total (high values of $N_{\mathrm{sig}}$ and $N_{90}>N_{\mathrm{sig}}$). Approximately 2/3 of the Auriga sample have $N_{\mathrm{sig}}>3$; all but two of the BJ05 sample have $N_{\mathrm{sig}}>5$. Furthermore, the BJ05 has no halos with $N_{\mathrm{sig}}<4$, and hence only limited overlap with the distribution of Aquarius and Auriga halos.
@@
This figure, for either measure of diversity, demonstrates that, when invoking the rule of thumb that ‘a few massive progenitors dominate a Milky Way-like halo’, ‘a few’ should be understood as ‘fewer than 10’, rather than ‘one or two’. The simulations we use are consistent with the earlier literature in this respect – for example, Deason et al. (2016), using an empirical method for assigning stellar masses to progenitors in $N$-body simulations of 45 halos with virial mass $10.2\times 10^{10}\,\mathrm{M_{\odot}}$, found that $\sim 50\,\%$ of their sample had $N_{\mathrm{prog}}>2$. Likewise, Monachesi et al. (2019) found a median $N_{90}\simeq 6.5$. On the basis of the models we examine here, a plausible estimate would be that the real Milky Way halo comprises dominant contributions from 2-5 progenitors of similar mass, and perhaps more, even if it is among the less diverse examples of a $10^{12}\,\mathrm{\mathrm{M_{\odot}}}$ system.
@@
Figure 4: Our diversity measure for the whole accreted stellar halo, $N_{\mathrm{sig}}$, compared to the alternative measure $N_{90}$ (e.g. Monachesi et al., 2019). Symbols correspond to model sets as shown in the legend, with labels showing the individual system numbers (Auriga) and letters (Aquarius). The color-coding of the square symbols for Aquarius corresponds to that used in other figures and in Cooper et al. (2010). The dashed line is the $1:1$ relation.
@@
Fig. 5 shows the full progenitor mass distribution of each halo, scaled in a way that isolates differences in the shape of these distributions. The horizontal axis shows the mass of each progenitor, expressed as a fraction of the total accreted mass. Progenitors are sorted along this axis, with the least massive on the left and the most massive on the right. The vertical axis shows the cumulative fraction of the total accreted mass provided by each progenitor. The figure therefore shows how a given halo is built up by the contributions of individual progenitors, in rank order by mass. Halos dominated by a single very massive progenitor (for example, Aquarius halo F) show a single large step; halos built up by many small contributions (typically the case for the BJ05 halos) show many small steps. The $N_{90}$ diversity measure is equivalent to the number of steps required to pass a line at $0.9$ on the vertical axis of this figure.
@@
Fig. 5 figure helps further to explain the differences between the diversity measures shown in Fig. 4. $N_{\mathrm{sig}}$ is sensitive to a small number of disproportionately massive progenitors, which Fig. 5 shows to be typical of the Auriga and Aquarius halos. In such cases, $N_{90}$ will give relatively more weight to smaller progenitors. In this paper, we use $N_{\mathrm{sig}}$ to quantify the diversity of accreted halos, because it more closely reflects the way in which observational estimates proceed in the Milky Way, starting from an approximate total mass, and weighted towards the most massive components, because those components are the most readily detected. $N_{\mathrm{sig}}$ is more robustly estimated from the incomplete data available in real observations; $N_{90}$ requires both an accurate estimate of the total accreted stellar halo mass and accurate accounting of the individual contributions from low-mass progenitors. Together with an estimate of the dominant progenitor mass from the peak metallicity of the halo stars, stellar population gradients may help to constrain $N_{90}$ (Monachesi et al., 2019); this approach works best where $N_{90}$ is small, and is likely to be more useful for external galaxies, where population gradients can be inferred from the color of the integrated light.

### v2

@@ 2.3 Auriga
Auriga is a suite of $N$-body magnetohydrodynamical zoom simulations of Milky Way–like galaxies from cosmological initial conditions (Marinacci & Pakmor, 2014; Grand et al., 2017, 2024). The simulations were carried out with the AREPO code (Springel, 2010). The original [^1] The recent Auriga data release described in Grand et al. (2024) includes nine additional halos with lower virial masses, $10^{11}--10^{12}\,\mathrm{\mathrm{M_{\odot}}}$. These are not considered here or in earlier papers about stellar halos in Auriga. Auriga suite comprises 30 halos simulated at a baseline resolution (level 4, L4), labeled Au01 to Au30. Of these, six were re-simulated at even higher resolution (level 3, L3). We use only the L4 halos in this work. Our study is concerned with the bulk structure of the halos, which for the most part is dominated by the most massive accreted progenitors. The lower resolution of L4 is therefore not as significant as the benefit of a much larger sample.
@@
In contrast to Aquarius and BJ05, Auriga includes star formation and the dynamical evolution of gas and stars self-consistently. This is particularly important for modeling the growth of the central galaxy and its response to individual accretion events. The distribution of orbits for halo stars should therefore better reflect both the phase space distribution of stars in satellites before they are disrupted, as well as interactions between the central galaxy and the accreted satellites (including stronger dynamical friction, more rapid disruption, and tidal shocking). In principle, a hydrodynamical treatment also allows for an ‘in situ halo’ formed by stars ‘kicked’ out of the disk during mergers or in gas clouds stripped from satellites, before they have time to mix into the gaseous halo or interstellar medium of the central galaxy (e.g. Cooper et al., 2015). However, as discussed by Cooper et al. (2017), it is not obvious that these effects will dominate the differences between particle tagging and hydrodynamical realizations of halos simulated from the same initial conditions: such differences are more likely to be dominated by model-to-model variations in the star formation histories of individual satellites predicted by different combinations of hydrodynamical and subgrid methods.
@@
The Auriga galaxies were chosen from a lower-resolution cosmological volume, based on a set of criteria to identify Milky Way analogs on the basis of their virial mass and isolation (see Grand et al., 2017, 2024). The resulting galaxies have radial scale lengths ranging from $2.16$ to $11.64\,\mathrm{kpc}$, most with prominent central bulge components. As described in Monachesi et al. (2019), their stellar halos, defined kinematically by stellar particles with circularity parameter $\epsilon<0.7$ and radius $>5$ kpc (to avoid the bulge region), have masses in the range $10^{9}$$10^{10}~{}\rm M_{\odot}$, with a significant contribution from stars formed in situ. The Auriga halos have a wide range of properties, consistent with the Aquarius halos in C10. Previous studies of Auriga have the above kinematic definition of the stellar halo, as well as other definitions; for consistency with the BJ05 and Aquarius models, we use a simple merger tree-based definition in this paper, described below.
@@ 2.4 Definition of the accreted stellar halo and its progenitors
Particles formed in long-lived satellites of the main halo may not be accreted into the stellar halo of the Milky Way analog until long after their parent halo has crossed the virial radius of the main branch (the infall time of the progenitor). A given progenitor ID therefore corresponds to a single infall time, but it may be associated with a wide range of individual particle accretion times. Satellite halos may continue to form stars after they become subhalos; if such stars are unbound from their parent satellite at $z=0$, we label them as accreted, consistent with the treatment of stars formed in the same progenitor before it became a subhalo of the main branch.
@@
Since we provide progenitor IDs as part of the datasets that accompany this paper, we remark on the following somewhat subtle detail. The progenitor ID indicates the branch from which a star was unbound when it became part of the stellar halo of the Milky Way analog: this is not necessarily the branch in which the star formed, because each subhalo joining the main branch has its own set of hierarchical progenitor branches. In the case of the most massive satellites, more than one of these ‘satellite of satellite’ branches may form stars. For example, say that galaxy B is a satellite of A, which itself becomes a satellite of the Milky Way analog. When this pair of galaxies crosses the virial radius of the Milky Way analog, B may be separated from A by tidal forces. Stars subsequently stripped from B will then be accreted directly into the stellar halo of the Milky Way analog. Our algorithm will assign a distinct progenitor ID to those stars (i.e., for those stars, B will be treated as a direct progenitor of the MW analog). However, stars from B that are first accreted by A (before or after A becomes a satellite of the MW analog) will instead be labeled with the progenitor ID of galaxy A. Consequently, it is possible for stars formed in one progenitor branch (the branch of galaxy B in this example) to have different progenitor IDs. This is consistent with our intended purpose here and is likely rare in practice [^3] Several technical subtleties are involved building merger trees by linking structures of halo-finding methods (like SUBFIND) across the snapshots of $N$-body simulations. For example, during complex merger events, the operational definition of the main branch as the most massive progenitor can be ambiguous, and links between ‘related’ branches or sections of branches can be broken (see Jiang et al., 2014, for extended discussion of the potential problems). The most significant impact of such uncertainties on our work is a small number of stellar particles that cannot be ascribed to any accretion event. We have not observed any cases in which these problematic particles contribute significant mass to the accreted stellar halo..
@@
In hydrodynamical simulations like Auriga, in situ stars can form directly on halo-like orbits or be scattered onto such orbits by gravitational interactions, after having formed in the central disk (e.g. Abadi et al., 2003; Zolotov et al., 2009; Cooper et al., 2015; Font et al., 2020). Alternative definitions of the stellar halo may count some fraction of these as ‘halo’ stars in order to make a more realistic comparison with observations (for example, selecting stars on very eccentric orbits). We prefer an idealized separation of accreted and in situ stars, using knowledge of the merger history from the simulations, because our purpose is mainly to compare different models (Auriga may overpredict the mass of in situ stars on halo orbits, which complicates any comparison that mixes accreted and in situ halo stars; Monachesi et al. 2019). In future work, it would be useful to examine how in situ stars affect inferences about the composition of the accreted stellar halo.
@@ 2.5.1 Stellar halo masses
Figure 1: Virial masses and accreted stellar halo masses of galaxies in the three model sets we use in this work (symbols, as shown in the legend). The star symbol shows recent estimates for the Milky Way (stellar halo mass $M_{\star}=1.4\pm 0.4\times 10^{9}\,\mathrm{M_{\odot}}$ (Deason et al., 2019, see also Mackereth & Bovy 2020); virial mass $M_{200}=0.97^{+0.24}_{-0.19}\times 10^{12}\,\mathrm{M_{\odot}}$ (Cautun et al. 2020; see also Wang et al. 2020)).
Figure 1 shows the virial mass and the total accreted stellar halo mass for each halo. The star symbol shows currently favored values for the Milky Way (Licquia & Newman, 2015; Wang et al., 2020). The models span slightly more than a factor of two in virial mass, with most systems more massive than the Milky Way. Accreted stellar halo masses range from $10^{8.5}$ to $10^{10.5}\,\mathrm{M_{\odot}}$ (in most cases, significantly larger than the Milky Way’s halo mass). The Auriga simulations account for most of this scatter, although this may be simply because of their larger sample size. However, there are some clear differences between the three sets of models. Auriga shows a weak trend in accreted stellar halo mass with virial mass. This is an expected consequence of the overall relationship between total stellar mass and virial mass (Purcell et al., 2007; Cooper et al., 2013). The typical stellar mass at fixed (peak) virial mass in Aquarius is somewhat less than that in Auriga (see Figure 2), leading to a lower total mass for the dominant progenitors of the accreted stellar halo in Aquarius. The apparently shallower trend of accreted stellar halo mass with virial mass in Aquarius is harder to interpret as a systematic difference with Auriga here, because of the small number of Aquarius halos and the large scatter among the Auriga galaxies. In the case of BJ05, the system-to-system scatter (which is due only to variations in star formation and accretion history at fixed virial mass) is clearly much smaller than that predicted by Auriga. This may be a consequence of the restricted range of accretion histories in BJ05, which excluded recent mergers with massive satellites.
As we discuss below, despite this large system-to-system scatter, the diversity of individual stellar halos shows clear systematic differences between the three model sets. We note that Auriga provides the simulations that are closest to measurements of the virial mass and stellar halo mass of the Milky Way (e.g. Au09, Au10, Au17, and Au22).
@@ 2.5.2 Progenitor masses
Figure 2: The relation between stellar mass and dark matter halo mass (prior to accretion) for progenitor satellites in the 6 Aquarius, 30 Auriga, and 11 BJ05 models. Progenitors from all individual halos in each simulation set are combined in this plot. The top and right-hand panels show the normalized distributions of halo mass and stellar mass, respectively.
Figure 2 shows the relationship between stellar mass and dark halo mass (approximately virial mass) for all progenitor satellites in Aquarius (blue), Auriga (pink), and BJ05 (gray). The total virial mass here is defined prior to infall of the progenitor to the main halo, hence this plot shows the ‘underlying’ relation, without the effects of tidal stripping on surviving satellites. The figure provides some insight into the different star formation efficiencies for satellites predicted by the models. They predict almost identical luminosity function shapes (right-hand histogram). Even well-resolved satellites in Auriga occupy somewhat less massive halos than satellites of comparable stellar mass in BJ05 or Aquarius (this can be seen in the shift of the halo mass distribution to lower halo mass for Auriga progenitors, as shown in the top histogram). The joint distributions show similar overall trends, but it is clear that Auriga progenitors have lower virial mass at fixed stellar mass (or equivalently, higher stellar mass at fixed virial mass), while BJ05 and Aquarius occupy almost the same region. The difference is greatest (up to $\sim 1$ order of magnitude) for the most massive satellites.
The most significant consequence of these differences is that the most massive progenitors in Auriga can contribute relatively more mass to the accreted stellar halo. Together with other effects such as dynamical friction, this is likely to lead to systematic differences between the diversity profiles of the models, as we discuss further below.
@@ 2.5.3 Overall density profiles
Figure 3 shows the overall distribution of accreted stars in the different models. The radial scales and outer power-law slopes are broadly consistent among the three models, but there are also clear systematic differences. The Auriga halos have the highest average density at all radii and are notably similar to one another in their outskirts, with $\lesssim 1~{}\mathrm{dex}$ variation in density between different galaxies at all radii $\gtrsim 10~{}\mathrm{kpc}$. The BJ05 halos have lower density than Auriga at radii $\lesssim 30~{}\mathrm{kpc}$ and even less galaxy-to-galaxy variation. The lower central densities of the BJ05 models (and perhaps also their lower average accreted stellar masses, at fixed virial mass) may be the result of the limitations of those models at higher redshift, as discussed above.
The Aquarius halos have lower density than most of the BJ05 and Auriga halos at radii $\gtrsim 10~{}\mathrm{kpc}$ and show greater variation from galaxy to galaxy, as well as a wider range of slopes and more pronounced inflections (‘breaks’). These breaks correspond to transitions between regions dominated by different combinations of progenitors, as described in C10. In some (not necessarily all) cases, the breaks correspond approximately to the apocenters of stars from individual dominant progenitors, and a pronounced break in the overall profile may indicate the mass of the accreted stellar halo overall is dominated by concentrated debris from a single progenitor (Deason et al., 2013). As we explore further below, the most significant difference between the Auriga and Aquarius stellar halos is that debris from individual massive satellites generally makes a dominant contribution to the overall density profile across a much wider range of radii in Auriga.
Figure 3: A comparison of the volume density profiles of the accreted stellar halos in each simulation. The six Aquarius halos are shown with individual colors. For simplicity, the 11 BJ05 halos and 30 Auriga halos are shown with gray and pink lines, respectively.
@@ 3.1 Definition of diversity
Other measures of overall stellar halo diversity (i.e., for the whole accreted stellar halo) have been used in the literature. For example, as noted above, BJ05 reported the fraction of accreted stellar mass contributed by the 15 most massive progenitors. Deason et al. (2016) reported the mass-weighted average progenitor stellar mass, which (as they explain) is directly related to our $N_{\mathrm{sig}}$. Another example is that of Monachesi et al. (2019), who used a measure we call $N_{90}$, defined as the number of the progenitors required to account for 90% of the accreted stellar halo mass, when accumulated in descending order of progenitor mass and rounded up to the nearest integer (a similar measure is discussed in C10). For example, a halo in which the most massive progenitor accounts for $80\%$ of the accreted stellar halo mass, and the second most massive progenitor for a further $15\%$, would have $N_{90}=2$. By construction, $N_{\mathrm{90}}\gtrsim N_{\mathrm{sig}}$. When a few massive progenitors account for most of the mass, but not more 90%, and the remaining mass up to 90% is dominated by a relatively large number of much less massive progenitors, then $N_{90}$ exceeds $N_{\mathrm{sig}}$ by a larger factor.
@@
Figure 4 compares $N_{90}$ with $N_{\mathrm{sig}}$ for the halos in the three simulation sets. It is clear that the different models predict different distributions of progenitor mass for the accreted stellar halo. The BJ05 and Auriga models contain examples of systems in which most of the accreted mass is contributed by many progenitors, each only a small fraction of the total (high values of $N_{\mathrm{sig}}$ and $N_{90}>N_{\mathrm{sig}}$). Approximately 2/3 of the Auriga sample have $N_{\mathrm{sig}}>3$; all but two of the BJ05 sample have $N_{\mathrm{sig}}>5$. Furthermore, the BJ05 has no halos with $N_{\mathrm{sig}}<4$, and hence only limited overlap with the distribution of Aquarius and Auriga halos.
@@
This figure, for either measure of diversity, demonstrates that, when invoking the rule of thumb that ‘a few massive progenitors dominate a Milky Way–like halo’, ‘a few’ should be understood as ‘fewer than 10’, rather than ‘one or two.’ The simulations we use are consistent with the earlier literature in this respect – for example, Deason et al. (2016), using an empirical method for assigning stellar masses to progenitors in $N$-body simulations of 45 halos with virial mass $10.2\times 10^{10}\,\mathrm{M_{\odot}}$, found that $\sim 50\%$ of their sample had $N_{\mathrm{prog}}>2$. Likewise, Monachesi et al. (2019) found a median $N_{90}\simeq 6.5$. On the basis of the models we examine here, a plausible estimate would be that the real Milky Way halo comprises dominant contributions from two to five progenitors of similar mass, and perhaps more, even if it is among the less diverse examples of a $10^{12}\,\mathrm{\mathrm{M_{\odot}}}$ system.
@@
Figure 4: Our diversity measure for the whole accreted stellar halo, $N_{\mathrm{sig}}$, compared to the alternative measure $N_{90}$ (e.g. Monachesi et al., 2019). Symbols correspond to model sets as shown in the legend, with labels showing the individual system numbers (Auriga) and letters (Aquarius). The color coding of the square symbols for Aquarius corresponds to that used in other figures and in Cooper et al. (2010). The dashed line is the $1:1$ relation.
@@
Figure 5 shows the full progenitor mass distribution of each halo, scaled in a way that isolates differences in the shape of these distributions. The horizontal axis shows the mass of each progenitor, expressed as a fraction of the total accreted mass. Progenitors are sorted along this axis, with the least massive on the left and the most massive on the right. The vertical axis shows the cumulative fraction of the total accreted mass provided by each progenitor. The figure therefore shows how a given halo is built up by the contributions of individual progenitors, in rank order by mass. Halos dominated by a single very massive progenitor (for example, Aquarius halo F) show a single large step; halos built up by many small contributions (typically the case for the BJ05 halos) show many small steps. The $N_{90}$ diversity measure is equivalent to the number of steps required to pass a line at $0.9$ on the vertical axis of this figure.
@@
Figure 5 figure helps further to explain the differences between the diversity measures shown in Figure 4. $N_{\mathrm{sig}}$ is sensitive to a small number of disproportionately massive progenitors, which Figure 5 shows to be typical of the Auriga and Aquarius halos. In such cases, $N_{90}$ will give relatively more weight to smaller progenitors. In this paper, we use $N_{\mathrm{sig}}$ to quantify the diversity of accreted halos, because it more closely reflects the way in which observational estimates proceed in the Milky Way, starting from an approximate total mass, and weighted toward the most massive components, because those components are the most readily detected. $N_{\mathrm{sig}}$ is more robustly estimated from the incomplete data available in real observations; $N_{90}$ requires both an accurate estimate of the total accreted stellar halo mass and accurate accounting of the individual contributions from low-mass progenitors. Together with an estimate of the dominant progenitor mass from the peak metallicity of the halo stars, stellar population gradients may help to constrain $N_{90}$ (Monachesi et al., 2019); this approach works best where $N_{90}$ is small, and is likely to be more useful for external galaxies, where population gradients can be inferred from the color of the integrated light.